# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ 4.3 An Example: k/G Inflation
To assess the stability of the model, we employ the methodology outlined in reference [15], imposing a small perturbation that deviates the field evolution from the analytic solution corresponding to a constant $\beta_{\text{\tiny{F}}}$. If the model is stable, it should be capable of returning to the constant-roll scenario after a small deviation. The evolution of the field and the constant-roll parameter $\beta_{\text{\tiny{F}}}$ under small perturbations in the field velocity are plotted in Figure 6. We observe that in the unperturbed constant-roll case, the field ultimately settles at the extrema of the effective potential, with the constant-roll parameter $\beta_{\text{\tiny{F}}}=-2.5$ remaining unchanged throughout. However, when the absolute value of the initial field velocity $|\dot{\phi}_{0}|$ is slightly reduced, the model promptly transitions to the corresponding dual attractor solution $\beta_{\text{\tiny{F}}}=-0.5$ (see (4.15)) and then settles at the extrema. The situation is a little different when we slightly increase $|\dot{\phi}_{0}|$. In this case, as the field reaches the extrema of the effective potential, where $\beta_{\text{\tiny{F}}}$ approaches $-3c_{s}^{2}$, it possesses sufficient kinetic energy to surpass this point. After surpassing the extrema, the field velocity gradually diminishes until it completely stops and reverses its direction. The model then immediately falls into the attractor dual solution $\beta_{\text{\tiny{F}}}=-0.5$ and eventually comes to rest at the extrema.
ptRefer to caption     Refer to caption
$(\phi-a_{1})\times 10^{2}$
@@
Figure 6: We have numerically integrated the field equation (4.4) for $\beta_{\text{\tiny{F}}}=-2.5$, $\gamma=1$, and $M=1$, to evaluate the time evolution of the field (left panel) and the constant-roll parameter $\beta_{\text{\tiny{F}}}=\ddot{\phi}/(H\dot{\phi})$ (right panel), in order to assess the stability of the constant-roll and propagation solution (4.31) under small perturbations. Initial values for the field and its velocity are taken as $-2$ and $-0.295-\delta$, respectively. Vertical grid lines correspond to the extremum of $\dot{\phi}$ in the $\delta=+0.01$ case. The horizontal grid line in the right panel shows the corresponding dual attractor solution $\beta_{\text{\tiny{F}}}=-0.5$.
@@
where for convenience we have define $M\equiv C_{3}$, $\bar{\beta}_{\text{\tiny{H}}}\equiv\sqrt[3]{-2\gamma\beta_{\text{\tiny{H}}}/9}$, and $\phi_{\text{\tiny{H}}}\equiv\bar{\beta}_{\text{\tiny{H}}}\phi$. For negative and positive $\beta_{\text{\tiny{H}}}$ the Hubble parameter (4.38) corresponds to the (2.13) and (2.14) cases, respectively. We have plotted this parameter for $\beta_{\text{\tiny{F}}}=-0.5$ with $\gamma=1$, as well as $\beta_{\text{\tiny{F}}}=0.5$ with $\gamma=-1$ in Figure 7 (top left panel). Likewise to the FCR case, these two scenarios share a common Hubble parameter but have different dynamical trends. In these specific choices, the parameter $\bar{\beta}_{\text{\tiny{H}}}$ is real and positive. This implies that for negative values of $\phi_{\text{\tiny{H}}}$, the Hubble parameter (4.38) adopts a sinusoidal form, while for positive $\phi_{\text{\tiny{H}}}$ values, it takes on a hyperbolic form. To pinpoint the critical points of the model, we introduce a new parameter $d_{n}$, defined as $d_{n}=-(n\pi/2)^{2/3}$. With this definition, zeros of Hubble parameter (4.38) are located at $\phi_{\text{\tiny{H}}}=d_{2s+1}$ and zeros of its derivative, $H_{,\phi}\propto\sqrt{\phi_{\text{\tiny{H}}}}\sinh\left(\phi_{\text{\tiny{H}}}^{3/2}\right)$, are located at $\phi_{\text{\tiny{H}}}=d_{2s}$, where $s$ is a non-negative integer. This section focuses on the model’s dynamics around $\phi_{\text{\tiny{H}}}=d_{0}=0$ and the first extrema points of the Hubble parameter and its derivative on the negative $\phi_{\text{\tiny{H}}}$ axis at $\phi_{\text{\tiny{H}}}=d_{2}\approx-2.15$ and $d_{1}\approx-1.35$, respectively.
ptRefer to caption    Refer to captionptRefer to caption     Refer to caption
@@
Figure 7: The plots of the Hubble parameter (4.38) (top left panel), number of $e$-folds (4.39) (top right panel), potential (4.40) and effective potential (4.42) (bottom left panel), and slow-roll parameters (4.44) and (4.45) (bottom right panel), once for $\beta_{\text{\tiny{F}}}=-0.5$ and $\gamma=1$ (solid lines) and once for $\beta_{\text{\tiny{F}}}=0.5$ and $\gamma=-1$ (dashed lines). The parameter $\mathcalboondox{V}_{0}$ in effective potential is taken to be $1.475$. Two vertical grid lines in all panels correspond to two $H_{,\phi}$ zeros at $\phi_{\text{\tiny{F}}}=d_{0}=0$ and $\phi_{\text{\tiny{F}}}=d_{2}\approx-2.15$.
@@
$N(\phi_{\text{\tiny{H}}})=\frac{1}{\beta_{\text{\tiny{F}}}}\ln\left(\left|\sinh\left(\phi_{\text{\tiny{H}}}^{3/2}\right)\right|\right)+N_{0}.$ (4.39)
The plot of this parameter is presented in Figure 7 (top right panel). In contrast to the FCR case, we find that while parameter $\mathcal{F}$ approaches zero at $\phi_{\text{\tiny{F}}}=0$, the number of $e$-folds exhibits a singularity at this point. This is because while the first term in equation (4.25) remains finite at this limit, the second term, $\ln\mathcal{F}$, diverges.
ptRefer to caption     Refer to caption
$N-N_{0}$
[…]
Figure 8: Another plot of the number of $e$-folds (4.39) (left panel), potential (4.40) (solid lines in right panel), and effective potential (4.42) (dashed lines in right panel) for $\gamma=1$ and $\mathcalboondox{V}_{0}=3.6$. Vertical grid line at $\phi_{\text{\tiny{F}}}=d_{2}\approx-2.15$ corresponds to singularities of $N(\phi_{\text{\tiny{H}}})$ functions.
Using (4.17) for the potential we get
@@
For the derivative of effective potential we use the equation (4.23) to obtain
@@
Integrating this equation, the effective potential is driven as
@@
Potential (4.40) and effective potential (4.42) are plotted in the bottom left panel of Figure 7. Extrema of effective potential are located at $d_{s}$, corresponding to zeros of the Hubble parameter and its derivative (see equation (4.23)). Comparing the effective potentials in Figures 4 and 7 around the extrema of the Hubble parameter ($a_{1}$ point in the FCR case and the $d_{2}$ point in the HCR case), we observe that the concavity of the effective potential in the HCR case resembles that of the FCR case. This is because, as we mentioned, at this critical point the derivatives of the effective potential in the FCR and PCR conditions are approximately equal, and the concavity of the effective potential is solely determined by the constant-roll parameter, regardless of the parameter $\mathcal{F}$. The concavity of the potential at $d_{2}$ point, however, depends on the sign of $\mathcal{F}$, which is why in both cases potential is concave. In the vicinity of $\phi_{\text{\tiny{H}}}=0$, where the sign of parameter $\mathcal{F}$ changes, the concavity of the potential undergoes a transition, but the concavity of the effective potential remains unchanged, again emphasizing that the effective potential is unaffected by the $\mathcal{F}$ function.
ptRefer to caption     Refer to caption
$(\phi-d_{2})\times 10^{2}$
@@
Figure 9: We have numerically integrated the field equation (4.4) for $\beta_{\text{\tiny{H}}}=-2.5$, $\gamma=1$, and $M=-1$, to evaluate the time evolution of the field (left panel) and the constant-roll parameter $\beta_{\text{\tiny{H}}}=\ddot{H}/(H\dot{H})$ (right panel), in order to assess the stability of the constant-roll and propagation solution (4.38) under small perturbations. Initial values for the field and its velocity are taken as $-2$ and $-1.36-\delta$, respectively. Vertical grid lines correspond to the extremum of $\dot{\phi}$ in the $\delta=+0.01$ case. The horizontal grid line in the right panel shows the corresponding dual attractor solution $\beta_{\text{\tiny{H}}}=-0.5$.
To derive the slow-roll parameters as a function of field, we can use equations (4.16). We will have:
@@
The bottom right panel of Figure 7 represents these parameters. As we can see, at critical points $\phi_{\text{\tiny{H}}}=d_{2s}$, the model exhibits the quasi-canonical behavior with $\epsilon_{1}$ approaching zero and $\epsilon_{2}\approx 2\beta_{\text{\tiny{H}}}$. When $\beta_{\text{\tiny{H}}}$ is negative, field moves toward extrema at $d_{2s}$, while for positive $\beta_{\text{\tiny{H}}}$ it moves away from these points. Therefore, similar to the FCR case, depending on the initial value of the field we will have different inflationary scenarios. In $\beta_{\text{\tiny{H}}}<0$ case, the field generally moves toward the extrema of the Hubble parameter (corresponding to singularities of the $N(\phi_{\text{\tiny{H}}})$) at $d_{2s}$ points, where we observe the quasi-canonical behavior, while in $\beta_{\text{\tiny{H}}}>0$ case, the field moves away from these points, gradually leaving the quasi-canonical phase and exits the inflation when $\epsilon_{1}$ reaches one. The exception is when both $\beta_{\text{\tiny{H}}}$ and initial value of the field are positive. In this case, field moves toward the positive infinity.
@@
To analyze the concavity of the potential and effective potential near $\beta_{\text{\tiny{H}}}=-3$, we have also plotted these parameters for $\beta_{\text{\tiny{H}}}=-4$, $-3$, and $-2$, in the vicinity of the critical point $\phi_{\text{\tiny{H}}}=d_{2}$ in Figure 8. A comparison of Figures 5 and 8 reveals that despite the general dissimilarity of potential and effective potential in the HCR case from the FCR, their concavity mirrors that of the FCR case. We also see that, unlike the FCR case, where the potential never maintained a USR profile, here we have a flat potential for $\beta_{\text{\tiny{H}}}=-3$, highlighting the effectiveness of the PCR condition for exploring the USR limit of a model.
@@
To check the stability of the model, we have numerically evaluated and plotted the dynamics of the field and the constant-roll parameter $\beta_{\text{\tiny{H}}}$ in response to slight perturbations in the field velocity in Figure 9. Similarly to the FCR case, we find that in the unperturbed case the constant-roll parameter $\beta_{\text{\tiny{H}}}$ retains a constant value $-2.5$ throughout the evolution, and upon imposing a tiny perturbation to the initial conditions, the model swiftly transitions into the corresponding dual attractor solution with $\beta_{\text{\tiny{H}}}=-2.5$, and the field eventually settles at the extremum of the effective potential at $\beta_{\text{\tiny{H}}}=-0.5$. Despite their shared features, it is essential to acknowledge that the FCR and HCR cases are not equivalent, as the transition trajectory is influenced by both the constant-roll condition and the parameter $\mathcal{F}$. For instance, as mentioned earlier, the value of the field velocity when the field passes the extrema of effective potential in the FCR case is independent of the form of the function $\mathcal{F}$, whereas it is dependent on $\mathcal{F}$ in the HCR case (see equations (4.11) and (4.23)).
5 Constant-Roll Kinetically Driven Inflation
In this section we discuss the $k$ inflation scenario [68, 69, 70, 71, 72, 73] as an example of models with varying sound speed that lead to three distinct solutions for the constant-roll conditions discussed in section 3. The equations of motion for the FCR and HCR conditions will be solved analytically to find two sets of constant-roll solutions. However, we lack the analyticity for the PCR condition, so we can only treat it numerically. For a better comparison between our three constant-roll conditions, we will introduce and analyze a concrete inflationary scenario consisting of multiple constant-roll stages where the power spectrum increases significantly and primordial black holes (PBHs) are formed. For each constant-roll condition, the power spectrum and other relevant parameters will be derived numerically and compared.
In the $k$ inflation we have
$P(X,\phi)=X+\frac{\gamma}{2}X^{2}-V(\phi)=\frac{\dot{\phi}^{2}}{2}\left(1+\frac{1}{4}\gamma\dot{\phi}^{2}\right)-V(\phi).$ (5.1)
Using this relation in equations (3.2), (3.3), and (3.4), for Fridmann equation and equations of motion we find
$3H^{2}=\frac{\dot{\phi}^{2}}{2}\left(1+\frac{3}{4}\gamma\dot{\phi}^{2}\right)+V,$ (5.2)
$\dot{H}=-\frac{\dot{\phi}^{2}}{2}\left(1+\frac{1}{2}\gamma\dot{\phi}^{2}\right),$ (5.3)
$\ddot{\phi}\left(1+\frac{3}{2}\gamma\dot{\phi}^{2}\right)+3H\dot{\phi}\left(1+\frac{1}{2}\gamma\dot{\phi}^{2}\right)+V_{,\phi}=0.$ (5.4)
Action of $k$ inflation (5.1) is clearly different from the constant propagation model (4.1). Thus, the model is expected to have a variable sound speed. Equation (3.5) can be used to determine sound speed as
$c_{s}^{2}=\frac{1+\gamma X}{1+3\gamma X}.$ (5.5)

### v3

@@ 4.3 An Example: k/G Inflation
To assess the stability of the model, we employ the methodology outlined in reference [21], imposing a small perturbation that deviates the field evolution from the analytic solution corresponding to a constant $\beta_{\text{\tiny{F}}}$. The evolution of the field and deviation of constant-roll parameter $\beta_{\text{\tiny{F}}}$ from the initially set value under small perturbations in the field velocity are plotted in Figure 6. We observe that in the unperturbed constant-roll case, the field ultimately settles at the extrema of the effective potential, with the constant-roll parameter $\beta_{\text{\tiny{F}}}=-2.5$ remaining unchanged throughout. However, when the absolute value of the initial field velocity $|\dot{\phi}_{0}|$ is slightly reduced, the model promptly transitions to the corresponding dual attractor solution $\beta_{\text{\tiny{F}}}=-0.5$ (see (4.15)) and then settles at the extrema. The situation is a little different when we slightly increase $|\dot{\phi}_{0}|$. In this case, as the field reaches the extrema of the effective potential, where $\beta_{\text{\tiny{F}}}$ approaches $-3$, it possesses sufficient kinetic energy to surpass this point. After surpassing the extrema, the field velocity gradually diminishes until it completely stops and reverses its direction. The model then immediately falls into the attractor dual solution $\beta_{\text{\tiny{F}}}=-0.5$ and eventually comes to rest at the extrema.
Refer to caption     Refer to caption
$(\phi-a_{1}/\bar{\beta}_{\text{\tiny{F}}})\times 10^{2}$
@@
Figure 6: We have numerically integrated the field equation (4.4) for $\beta_{\text{\tiny{F}}}=-2.5$, $\gamma=1$, and $M=1$, to evaluate the time evolution of the field (left panel) and the constant-roll parameter $\beta_{\text{\tiny{F}}}=\ddot{\phi}/(H\dot{\phi})$ (right panel), in order to assess the stability of the constant-roll and propagation model (4.31) under small perturbations of field velocity, $\delta$. Initial values for the field and its velocity are taken as $-2$ and $-0.295-\delta$, respectively. Vertical grid lines correspond to the time when field stops and reverses its direction in the $\delta=+0.01$ case. The horizontal grid line in the right panel shows the corresponding dual attractor solution $\beta_{\text{\tiny{F}}}=-0.5$.
@@
where for convenience we have defined $M\equiv C_{3}$, $\bar{\beta}_{\text{\tiny{H}}}\equiv\sqrt[3]{-2\gamma\beta_{\text{\tiny{H}}}/9}$, and $\phi_{\text{\tiny{H}}}\equiv\bar{\beta}_{\text{\tiny{H}}}\phi$. For negative and positive $\beta_{\text{\tiny{H}}}$ the Hubble parameter (4.38) corresponds to the (2.13) and (2.14) cases, respectively. We have plotted this parameter for $\beta_{\text{\tiny{H}}}=-0.5$ with $\gamma=1$, as well as $\beta_{\text{\tiny{H}}}=0.5$ with $\gamma=-1$ in Figure 7 (top left panel). Similar to the FCR case, these two scenarios share a common Hubble parameter but have different dynamical trends. In these specific choices, the parameter $\bar{\beta}_{\text{\tiny{H}}}$ is real and positive. This implies that for negative values of $\phi_{\text{\tiny{H}}}$, the Hubble parameter (4.38) adopts a sinusoidal form, while for positive $\phi_{\text{\tiny{H}}}$ values, it takes on a hyperbolic form. To pinpoint the critical points of the model, we introduce a new parameter $d_{n}$, defined as $d_{n}=-(n\pi/2)^{2/3}$. With this definition, zeros of Hubble parameter (4.38) are located at $\phi_{\text{\tiny{H}}}=d_{2s+1}$ and zeros of its derivative, $H_{,\phi}\propto\sqrt{\phi_{\text{\tiny{H}}}}\sinh(\phi_{\text{\tiny{H}}}^{3/2})$, are located at $\phi_{\text{\tiny{H}}}=d_{2s}$, where $s$ is a non-negative integer. This section focuses on the model’s dynamics around the first peak of the Hubble parameter on the negative $\phi_{\text{\tiny{H}}}$ axis at $\phi_{\text{\tiny{H}}}=d_{4}$, and we restrict our discussion to the $d_{5}<\phi_{\text{\tiny{H}}}<d_{3}$ interval where Hubble parameter is positive for $M>0$.
Refer to caption    Refer to captionRefer to caption     Refer to caption
@@
Figure 7: The plots of the Hubble parameter (4.38) (top left panel), number of $e$-folds (4.39) (top right panel), potential (4.40) and effective potential (4.42) (bottom left panel), and slow-roll parameters (4.44) and (4.45) (bottom right panel), once for $\beta_{\text{\tiny{H}}}=-0.5$ and $\gamma=1$ (solid lines) and once for $\beta_{\text{\tiny{H}}}=0.5$ and $\gamma=-1$ (dashed lines). The parameter $\mathcalboondox{V}_{0}$ in effective potential is taken to be $1.5$. Vertical grid lines in all panels correspond to the extrema of $H_{,\phi}$ at $\phi_{\text{\tiny{H}}}=d_{4}\approx-3.4$.
@@
$N(\phi_{\text{\tiny{H}}})=\frac{1}{\beta_{\text{\tiny{H}}}}\ln\left(\left|\sinh\left(\phi_{\text{\tiny{H}}}^{3/2}\right)\right|\right)+N_{0}.$ (4.39)
This function is shown in Figure 7 (top right panel). Using (4.17) for the potential we get
@@
For the derivative of effective potential equation (4.23) gives
@@
Integrating this equation, the effective potential is derived as
@@
Potential (4.40) and effective potential (4.42) are plotted in the bottom left panel of Figure 7. Zeros of the Hubble parameter and its derivative coincide with extrema of effective potential (see equation (4.23)). Comparing the behavior of effective potential in Figures 4 and 7 around the extrema of the Hubble parameter ($a_{1}$ point in the FCR case and $d_{4}$ point in the HCR case), we observe that the concavity of both models are alike. This is because, as we mentioned, at this critical point the derivatives of the effective potential in the FCR and PCR conditions are approximately equal, and the concavity of the effective potential is solely determined by the constant-roll parameter, regardless of the parameter $\mathcal{F}$. The concavity of the potential at $d_{4}$ point, however, depends on the sign of $\mathcal{F}$, which is why in both cases potential is concave.
Refer to caption     Refer to caption
$N-N_{0}$
[…]
Figure 8: Another plot of the number of $e$-folds (4.39) (left panel), potential (4.40) (solid lines in right panel), and effective potential (4.42) (dashed lines in right panel) for $\gamma=1$ and $\mathcalboondox{V}_{0}=3.382$. Vertical grid line at $\phi_{\text{\tiny{H}}}=d_{4}\approx-3.4$ corresponds to the singularity of function $N(\phi_{\text{\tiny{H}}})$.
Refer to caption     Refer to caption
$(\phi-d_{4}/\bar{\beta}_{\text{\tiny{H}}})\times 10^{2}$
@@
Figure 9: We have numerically integrated the field equation (4.4) for $\beta_{\text{\tiny{H}}}=-2.5$, $\gamma=1$, and $M=1$, to evaluate the time evolution of the field (left panel) and the constant-roll parameter $\beta_{\text{\tiny{H}}}=\ddot{H}/(2H\dot{H})$ (right panel), in order to assess the stability of the constant-roll and propagation model (4.38) under small perturbations of field velocity, $\delta$. Initial values for the field and its velocity are taken as $-2$ and $-1.358-\delta$, respectively. Vertical grid lines correspond to the time when field stops and reverses its direction in the $\delta=+0.01$ case. The horizontal grid line in the right panel shows the corresponding dual attractor solution $\beta_{\text{\tiny{H}}}=-0.5$.
To derive the slow-roll parameters as a function of field, one can use equation (4.16) to obtain
@@
The bottom right panel of Figure 7 represents these parameters. As we can see, at critical point $\phi_{\text{\tiny{H}}}=d_{4}$, the model exhibits the quasi-canonical behavior with $\epsilon_{1}$ approaching zero and $\epsilon_{2}\approx 2\beta_{\text{\tiny{H}}}$. One can think of different inflationary scenarios depending on the initial value of the field. In $\beta_{\text{\tiny{H}}}<0$ case, the field generally moves toward the extrema of the Hubble parameter (corresponding to singularities of $N(\phi_{\text{\tiny{H}}})$) at $d_{4}$ point, where we observe the quasi-canonical behavior, while in $\beta_{\text{\tiny{H}}}>0$ case, the field moves away from these points, gradually leaving the quasi-canonical phase and exits the inflation when $\epsilon_{1}$ reaches one. The exception is when both $\beta_{\text{\tiny{H}}}$ and initial value of the field are positive. In this case, field moves toward the positive infinity.
@@
To analyse the concavity of the potential and effective potential near $\beta_{\text{\tiny{H}}}=-3$, we have also plotted these parameters for $\beta_{\text{\tiny{H}}}=-4$, $-3$, and $-2$, in the vicinity of the critical point $\phi_{\text{\tiny{H}}}=d_{4}$ in Figure 8. A comparison of Figures 5 and 8 reveals that despite the general dissimilarity of the potential and effective potential in the HCR case from the FCR, their concavity mirrors that of the FCR case. We also see that, unlike the FCR case, where the potential does not support the USR profile, here we have a flat potential for $\beta_{\text{\tiny{H}}}=-3$, highlighting the effectiveness of the PCR condition for exploring the USR limit of a model.
@@
To check the stability of the model, we have numerically evaluated and plotted the dynamics of the field and the constant-roll parameter $\beta_{\text{\tiny{H}}}$ in response to slight perturbations in the field velocity in Figure 9. Similarly to the FCR case, we find that in the unperturbed case the constant-roll parameter $\beta_{\text{\tiny{H}}}$ retains a constant value $-2.5$ throughout the evolution, and upon imposing a tiny perturbation to the initial conditions, the model swiftly transitions into the corresponding dual attractor solution with $\beta_{\text{\tiny{H}}}=-0.5$, and the field eventually settles at the extremum of the effective potential at $\phi_{\text{\tiny{H}}}=d_{4}$. Despite their shared features, it is essential to acknowledge that the FCR and HCR cases are not equivalent, as the transition trajectory is influenced by both the constant-roll condition and the function $\mathcal{F}$. For instance, as mentioned earlier, the value of the field velocity when the field passes the extrema of effective potential in the FCR case is independent of the form of the function $\mathcal{F}$, whereas it depends on $\mathcal{F}$ in the HCR case (see equations (4.11) and (4.23)).
5 Constant-Roll Kinetically Driven Inflation and Scalar Perturbations
In this section, we will examine another example of non-canonical constant-roll models, in which sound speed is not constant and $Q_{,\phi}=0$. The equations of motion for the three different conditions discussed in section 3 will be solved. For a better comparison between dynamics of the model under these conditions, we will introduce and analyse a concrete inflationary scenario consisting of multiple constant-roll stages. We will discuss the background in detail and study the perturbations. In this scenario, there are phases with non-slow-roll evolution able to amplify the size of curvature perturbation to the level required for triggering primordial black holes (PBHs).
5.1 Background Dynamics
We consider a special case of the kinetically driven inflation (or $k$ inflation) as [69, 70, 71, 72, 73, 74]
$P(X,\phi)=X+\frac{\gamma}{2}X^{2}-V(\phi),$ (5.1)
where $\gamma$ is a constant with dimension of $M^{-4}$. Using this relation in equations (3.2)-(3.4), for the Friedmann and field equations we get
$6H^{2}=X\left(2+3\gamma X\right)+2V,$ (5.2)
$\dot{H}+X\left(1+\gamma X\right)=0,$ (5.3)
$\ddot{\phi}+3c_{s}^{2}H\dot{\phi}+\mathcalboondox{V}_{,\phi}=0,$ (5.4)
where the sound speed is given by
$c_{s}^{2}=\frac{1+\gamma X}{1+3\gamma X},$ (5.5)
and the derivative of effective potential (3.6) is related to potential as
$\mathcalboondox{V}_{,\phi}=\frac{V_{,\phi}}{1+3\gamma X}.$ (5.6)
Furthermore, the two first slow-roll parameters are expressed as
$\epsilon_{1}=\frac{X(1+\gamma X)}{H^{2}},\quad\quad\epsilon_{2}=2\lambda\frac{\ddot{\phi}}{H\dot{\phi}}.$ (5.7)
We will also need the time rate of sound speed variation (A.4), which is given by
$q_{1}=\frac{\left(1-c_{s}^{2}\right)\left(1-3c_{s}^{2}\right)}{2c_{s}^{2}}\frac{\ddot{\phi}}{H\dot{\phi}}.$ (5.8)